OpenStudy (anonymous):

Find the measure of an interior angle an exterior angle of a regular polygon with 5 sides .

OpenStudy (anonymous):

to find the interior angle, it is (n-2)180 so it would be 5-2(180) which is 3*180

OpenStudy (p0sitr0n):

360/5=72 (180-72) = 108 - internal 360-internal = external

OpenStudy (anonymous):

that is the sum of internal angles. to get one interior angle, divide the sum of interior angles by 5.

OpenStudy (anonymous):

To find all exterior angles, it is alway 360 in a regular polygon. to find 1 exterior angle, it would be 360/5

OpenStudy (anonymous):

so: 1 interior angle= 540/5 1 exterior angle=360/5 hope this helped! :)